OpenStudy (kayders1997):

Find the integral of x(2x+5)^8dx

OpenStudy (kayders1997):

@satellite73

OpenStudy (kayders1997):

@ganeshie8

OpenStudy (vishweshshrimali5):

\(\int x(2x+5)^8 dx\) Put, 2x + 5 = u

OpenStudy (vishweshshrimali5):

What will be x and dx in terms of u ?

OpenStudy (kayders1997):

Du=2dx

OpenStudy (kayders1997):

I don't know if that's right though because you need to get back to xdx

OpenStudy (anonymous):

you can use \[u=2x+5\]then solve the equation for \(x\) in terms of \(yu\)

OpenStudy (anonymous):

i meant solve for \(x\) in terms of \(u\)

OpenStudy (vishweshshrimali5):

du = 2dx is correct What about x in terms of u ?

OpenStudy (kayders1997):

But wouldn't that be du=2dx is there something I could do different?

OpenStudy (kayders1997):

Okay

OpenStudy (anonymous):

it is kind of going to suck, but you can do it

OpenStudy (vishweshshrimali5):

Yeps :P

OpenStudy (anonymous):

\[x=\frac{u-5}{2}\]

OpenStudy (kayders1997):

I've never done a problem like this

OpenStudy (anonymous):

maybe it is a typo

OpenStudy (kayders1997):

It's out of a calculus book

OpenStudy (vishweshshrimali5):

No problem then.. Your integral will become: \cfrac{1}{2}\[\int \cfrac{u - 5}{2}u^5 du\]

OpenStudy (anonymous):

ok then grind it till you find it btw calc books have typos too lots of them

OpenStudy (kayders1997):

Why u^5? Hmmmm

OpenStudy (vishweshshrimali5):

Sorry the integral is: \[\cfrac{1}{2}\int \cfrac{u - 5}{2}u^8 du\] \[=\cfrac{1}{4} \int (u^9 - 5u^8) du\]

OpenStudy (vishweshshrimali5):

Can you solve this?

OpenStudy (kayders1997):

Yeah just a second I'm processing the information you just gave

OpenStudy (vishweshshrimali5):

Sure

OpenStudy (kayders1997):

The one thing I don't understand us why 1/2 is in front of the integral?

OpenStudy (vishweshshrimali5):

Recall that we got du = 2dx => dx = du/2 So, this is where we got that that 1/2

OpenStudy (kayders1997):

Oh okay!

OpenStudy (kayders1997):

I got 1/40(2x+5)^10-5/36(2x+5)^9

OpenStudy (kayders1997):

Sorry if that's hard to read

OpenStudy (vishweshshrimali5):

Perfect!! Don't forget the integration constant C

OpenStudy (kayders1997):

I always forget it!!!! Ugh!!!!

OpenStudy (vishweshshrimali5):

:) It's alright!

OpenStudy (kayders1997):

Thank you so much :)

OpenStudy (vishweshshrimali5):

Glad to be of help :)

OpenStudy (kayders1997):

Definitely a fan

OpenStudy (kayders1997):

Would you be able to help with a different topic soon?

OpenStudy (kayders1997):

I'm not good at intragration by parts